OpenStudy (anonymous):

please help me out with this one- a right circular cone has a surface area of 718 square inches. what dimensions (radius and height) will result in a maximum volume?

OpenStudy (anonymous):

I began the problem, and now it after 30 mins of work it is a nightmare.

OpenStudy (mertsj):

What class is this for--calculus?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so far, i've taken the SA eqn and tried to optimize it with substitution, but then after finding a derivative it just is bad.

OpenStudy (mertsj):

If you could find satellite of myininaya they could help you

OpenStudy (amistre64):

this one, jwg lol

OpenStudy (mertsj):

yep

OpenStudy (amistre64):

is there a formula for the surface area of a rt circular cone?

OpenStudy (mertsj):

SA=pir^2+pirl

OpenStudy (amistre64):

and perferably one for volume as well to relate the 2?

OpenStudy (amistre64):

1/3 something is all i can remember

OpenStudy (amistre64):

1/3 volume of a cylinder perhaps?

OpenStudy (mertsj):

V=-1/3pir^2h

OpenStudy (mertsj):

Exactly

OpenStudy (amistre64):

ok; looks like we are given an value for SA so that we can find an r perhaps?

OpenStudy (mertsj):

The problem is the slant height adds a variable.

OpenStudy (amistre64):

i take it thats the "l" part in the SA

OpenStudy (mertsj):

yes

OpenStudy (amistre64):



OpenStudy (mertsj):

precisely

OpenStudy (amistre64):

l = sqrt(h^2 + r^2) then 718 = pi r^2 + pi r(sqrt(h^2+r^2)) 718/pir = r + (sqrt(h^2+r^2)) (718/pir) - r = sqrt(h^2+r^2) [(718/pir) - r]^2 = h^2+r^2 [(718/pir) - r]^2 - r^2 = h^2 i wonder of thats going according to plan :)

OpenStudy (amistre64):

at any rate we can define h in terms of r; aint pretty but doable

OpenStudy (mertsj):

That should be progress. Wonder where the asker is.

OpenStudy (amistre64):

fetal position, sobbing maybe lol

OpenStudy (amistre64):

V = 1/3 pi r^2 h \[V=\frac{pi\ r^2\ \sqrt{(\frac{718}{pi\ r}-r)^2-r^2}}{3}\] looks about right?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

Now I suppose we have to differentiate that and set to 0?

OpenStudy (amistre64):

the 1/3 pi is a constant that can be put to the side and we are left with deriving:\[V=r^2\ ((\frac{718}{pi\ r}-r)^2-r^2)^{1/2}\]

OpenStudy (amistre64):

yep, find the zeros and the undefineds to test

OpenStudy (amistre64):

\[V'=r'^2\ ((\frac{718}{pi\ r}-r)^2-r^2)^{1/2}+r^2\ ((\frac{718}{pi\ r}-r)^2-r^2)'^{1/2}\] \[V'=2r\ ((\frac{718}{pi\ r}-r)^2-r^2)^{1/2}+\frac{1}{2}r^2\ ((\frac{718}{pi\ r}-r)^2-r^2)^{-1/2}*((\frac{718}{pi\ r}-r)^2-r^2)'\] \[V'=2r\ ((\frac{718}{pi\ r}-r)^2-r^2)^{1/2}+\frac{1}{2}r^2\ ((\frac{718}{pi\ r}-r)^2-r^2)^{-1/2}*(\frac{718}{pi\ r}-r)'^2-r'^2\] \[V'=2r\ ((\frac{718}{pi\ r}-r)^2-r^2)^{1/2}+\frac{1}{2}r^2\ ((\frac{718}{pi\ r}-r)^2-r^2)^{-1/2}*2(\frac{718}{pi\ r}-r)(\frac{718}{pi\ r}-r)'-2r\] yep, its a doozie; the wolf might be quicker

OpenStudy (mertsj):

That's what I did.

OpenStudy (amistre64):

..... slacker!! lol

OpenStudy (mertsj):

proudly so

OpenStudy (mertsj):

r=7.56

OpenStudy (amistre64):

yeah, i was gonna say around 8

OpenStudy (amistre64):

plug that value in to get the V max

OpenStudy (mertsj):

I just ignored the miserable denominator and set the numerator to 0.

OpenStudy (mertsj):

Must first find the height

OpenStudy (amistre64):

[(718/pir) - r]^2 - r^2 = h^2 plug in yer r

OpenStudy (mertsj):

yep

OpenStudy (amistre64):

i dunno if solving in terms of h would have been easier

OpenStudy (mertsj):

tooooooooooooooooooooooooooooo late

OpenStudy (amistre64):

lol, since h was buried i assume not

OpenStudy (mertsj):

h= 21.37

OpenStudy (amistre64):

that should do it then

OpenStudy (mertsj):

Except for checking to make sure those values result in the correct surface area.

OpenStudy (amistre64):

it has to, we did everything correctly

OpenStudy (mertsj):

Of course...no mistakes...ever!!

OpenStudy (amistre64):

they wanted max volume

OpenStudy (amistre64):

the wolf helped so we can always pass the blame ;)

OpenStudy (mertsj):

working on that.

OpenStudy (mertsj):

1279.02= max volume

OpenStudy (mertsj):

Good to have a scapegoat.

OpenStudy (mertsj):

Hey!! It's right. SA = 717.9758

OpenStudy (mertsj):

Thank you very much.